A gray kangaroo can bound across level ground with each jump carrying it 8.9 m from the takeoff point. Typically the kangaroo leaves the ground at a 21 ∘ angle. If this is so:

A-) What is its takeoff speed?

B-) What is its maximum height above the ground?
Answer #1

A)

The motion of kangaroo will be projectile motion

Range of a projectile motion R = u^2 sin(2 * theta) / g

8.9 = u^2 * sin(2 * 21) / 9.81

=> u^2 = 130.48

=> u = 11.423 m / sec

B)

Maximum height in projectile motion (h) = u^2 * sin^2(theta) / (2 * g)

=> h = (11.423)^2 * sin^2(21) / (2 * 9.81)

=> h = 6.651 * sin^2(21) = 0.854 meters